# Code Prompting Elicits Conditional Reasoning Abilities in Text+Code LLMs

### v2

@@ Abstract
Reasoning is a fundamental component of language understanding. Recent prompting techniques, such as chain of thought, have consistently improved LLMs’ performance on various reasoning tasks. Nevertheless, there is still little understanding of what triggers reasoning abilities in LLMs in the inference stage. In this paper, we introduce code prompting, a chain of prompts that transforms a natural language problem into code and directly prompts the LLM using the generated code without resorting to external code execution. We hypothesize that code prompts can elicit certain reasoning capabilities of LLMs trained on text and code and utilize the proposed method to improve conditional reasoning, the ability to infer different conclusions depending on the fulfillment of certain conditions. We find that code prompting exhibits a high-performance boost for multiple LLMs (up to percentage points on GPT 3.5, on Mixtral, and on Mistral) across multiple conditional reasoning datasets. We then conduct comprehensive experiments to understand how code prompts trigger reasoning abilities and which capabilities are elicited in the underlying models. Our analysis of GPT 3.5 reveals that the code formatting of the input problem is essential for performance improvement. Furthermore, code prompts improve sample efficiency of in-context learning and facilitate state tracking of variables or entities.111Our code and prompts are available at this URL.
@@
Haritz Puerto, Martin Tutek, Somak Aditya, Xiaodan Zhu, Iryna Gurevych Ubiquitous Knowledge Processing Lab (UKP Lab), TU Darmstadt and Hessian Center for AI (hessian.AI) IIT Kharagpur, Queen’s University https://www.ukp.tu-darmstadt.de
@@ 1 Introduction
Reasoning is a fundamental component of both human and artificial intelligence (AI) and the backbone of many NLP tasks, such as question answering and textual entailment. Recently, intensive studies have been performed on mathematical (Patel et al., 2021; Chen et al., 2021; Cobbe et al., 2021), logical (Liu et al., 2020, 2023a; Sinha et al., 2019), and commonsense reasoning (Madaan et al., 2022; Liu et al., 2022a, b; Wang et al., 2023). Conditional reasoning, a primary yet complex reasoning ability that draws alternative conclusions depending on the fulfillment of certain conditions, remains understudied. These conditions are stated in the text, making the problem self-contained, which allows us to study the semantic inferencing capabilities of the underlying model, i.e., identifying relevant premises and ascertaining the presence of total evidence Nolt et al. (1988); Cabria and Magnini (2014) without the requirement for, and confounding effects of external knowledge. Conditional reasoning is also a fundamental form of logical reasoning useful in many practical scenarios, such as answering real-world questions about the eligibility for a visa or a loan. Despite the recent introduction of some benchmarks (Saeidi et al., 2018; Sun et al., 2022; Kazemi et al., 2023), conditional reasoning abilities of LLMs remain unknown.
@@
Recently, researchers have improved performance on reasoning tasks by combining LLMs with symbolic interpreters such as the Python runtime (Gao et al., 2023; Chen et al., 2023; Lyu et al., 2023) or SATisfiability solvers (Ye et al., 2023). Here, LLMs pretrained on code or a combination of text and code (henceforth text+code LLM) are used to convert the input task into a symbolic language (e.g., Python or SAT problems), which is then fed into an external interpreter to make use of its symbolic execution. In such a setup, the LLM is mainly used for the linguistic dimension of reasoning (i.e., identifying the premises from text) and planning how to solve the problem, while the logical reasoning component (i.e., solving the actual logical problem) is offloaded to an external execution module, confounding our understanding of the reasoning abilities of the underlying model. In particular, the fundamental questions of what contributes to the reasoning abilities and how reasoning abilities are triggered remain open. Nevertheless, pretraining on code is considered an important component that contributes to and explains the improved reasoning ability of LLMs. State-of-the-art LLMs such as GPT 3.5 (Kojima et al., 2022), GPT 4 (OpenAI, 2023), Mixtral (Jiang et al., 2024), and Mistral 7B (Jiang et al., 2023) have been pretrained on both text and code and have demonstrated considerable boosts in many reasoning benchmarks.
In this work, we focus on improving the logical reasoning dimension of semantic inference and investigate which aspects of code prompts, and in which way, elicit conditional reasoning abilities of several text+code LLMs (GPT 3.5, Mixtral, and Mistral), across three datasets: (1) ConditionalQA (Sun et al., 2022), (2) BoardgameQA (Kazemi et al., 2023) and (3) ShARC (Saeidi et al., 2018). To understand the benefit of code as an intermediate representation, we devise a chain of prompts, code prompting, that transform a natural language (NL) task into code and directly prompts the LLM with the generated code. The code contains the logical structure needed to solve the problem, along with the original natural language text as code comments. An illustration is provided in Figure 1. This setup has multiple benefits. Firstly, we fully utilize the LLM’s reasoning abilities without offloading any subtask to an external interpreter. In this way, we can conduct a fair comparison between text and code prompts without external variables such as interpreters. Secondly, enforcing compilability and executability of the generated code is not required, resulting in fewer interpreter-driven errors.
Our contributions are summarized as follows:
We devise a chain of prompts that transforms a NL task into code to trigger conditional reasoning abilities in text+code LLMs.
@@
We conduct an extensive analysis to understand why code prompts efficiently elicit conditional reasoning abilities, and show that prompting with code results in largely improved variable state tracking.
@@ 2 Background and Related Work
Semantic Inference.
It is important to understand the framework within which we conduct our experiments. Question answering can be seen as a chain of semantic inferences. Semantic inference can be decomposed into (1) the linguistic dimension, which identifies implications from natural language text and (2) the logical dimension, where reasoning, deductive, inductive or abductive, is conducted given the identified premises (Cabria and Magnini, 2014). The logical dimension resides on four major criteria which a reasoning model should validate Nolt et al. (1988): (a) the truth of premises, (b) validity and inductive probability of premises, (c) relevance of premises to the conclusion, and (d) requirement of total evidence. In the scope of our work, we focus on improving deductive reasoning along the logical dimension. We assume the truthfulness of information presented within the datasets, making the task of the analyzed LLM to (c) identify relevant premises and (d) ascertain whether total required evidence is present, which is the focus of conditional reasoning.
Code LLMs.
@@
Most works that generate code to solve natural language tasks use an external symbolic interpreter to run the resulting code. Chen et al. (2023) and Gao et al. (2023) showed consistent gains on mathematical problems, symbolic reasoning, and algorithmic problems by using LLMs aided by external code interpreters. Lyu et al. (2023) further showed the improvement in multi-hop QA, planning, and relational inference. In contrast, Ye et al. (2023) used an external automated theorem prover with declarative code and showed consistent gains w.r.t. imperative code-interpreter-aided LLMs on arithmetic reasoning, logical reasoning, symbolic reasoning, and regex synthesis tasks. Lastly, Pan et al. (2023) did not use any interpreter and instead created programs composed of multiple subroutines and used smaller specialized models to run them. In this way, they outperform text prompts on text LLMs for fact-checking tasks.
All these works employ an external solver system to run the code; therefore, the LLM is not conducting part of the reasoning. Some works (Madaan et al., 2022; Liu et al., 2023b) suggest that LLMs trained on code may possess superior reasoning abilities than LLMs trained only on natural language text. Madaan et al. (2022) observed improved commonsense reasoning, while Liu et al. (2023b) report superior causal reasoning. The latter work is based on the intuition that large amounts of if statements in the pretraining corpus enhance causal reasoning abilities because they represent explicit causal relations. They conducted experiments on abductive and counterfactual reasoning tasks and showed that translating NL problems into code and then generating functions that return the answer to the problem with a code LLM outperforms prompting the same NL task in a text LLM. However, most popular LLMs are trained on text and code (e.g., GPT 3.5; Kojima et al. 2022, GPT 4; OpenAI 2023). Therefore, it remains unclear whether code prompts also outperform text prompts in text+code models and, if so, why code prompts elicit reasoning abilities. In our work, we aim to answer these questions.
To the best of our knowledge, only the concurrent work of Hussain et al. (2023) investigates the conditional reasoning abilities of LLMs. However, they only analyze the abilities of text-only LLMs after training them on ConditionalQA (Sun et al., 2022). Instead, we investigate how to use prompts to elicit conditional reasoning in text+code LLMs.
@@ 3 Code Prompting
3.1 Definition
We hypothesize that querying text+code LLMs with instances translated to code will result in improved conditional reasoning capabilities. Our hypothesis is motivated by prior works showing that program-aided LMs exhibit superior results on reasoning tasks than regular text prompts (Gao et al., 2023; Chen et al., 2023; Lyu et al., 2023; Ye et al., 2023). Thus, it comes naturally that prompting text+code LLMs with instances translated to code could cause the underlying model to exhibit superior reasoning abilities when compared to text prompts. We formalize our hypothesis as follows:
where is a problem instance from the set of problems , is an evaluation function that returns the quality of an answer for a given problem, and and are functions that create a natural language Text prompt and Code prompt, respectively, for the same given problem.
We define code prompts as prompts that model a natural language (NL) problem with code. The code contains the logical structure needed to solve the problem, along with the original natural language text as code comments. To solve an NL task with code prompts, we define a chain of prompts that i) transform the NL text into code, and ii) use this code to generate the answer in natural language. Figure 1 illustrates this pipeline.
The prompt transforming the NL problem into code is composed of code that closely follows the original NL text. In particular, it creates variables for key entities in the question and documents and if blocks for each conditional statement in the documents. Figure 2 exemplifies this transformation.
3.2 Coding Features
@@ 4.1 Datasets
We focus on the QA task of CondQA. For BGQA, we focus on the main partition, which includes three subsets BGQA-1, BGQA-2, and BGQA-3, where the number indicates the reasoning hops needed to answer. Lastly, while ShARC encompasses dialogue generation, we aim to evaluate specific capabilities unrelated to conversational flow. Therefore, we isolated the QA pairs from the provided dialogues, resulting in a dataset where the model has to answer yes, no, or not enough information.222In the full task, not enough information would trigger another step in a pipeline to generate a follow-up question.
We include more details about the datasets in Appendix A, a formal definition of the prompts in Appendix B, and examples in Appendix K.
4.2 LLM Setup
We perform our study using gpt-35-turbo333https://platform.openai.com/docs/models/gpt-3-5-turbo through the Azure OpenAI service, Mixtral 7x8B (Jiang et al., 2024), and Mistral 7B (Jiang et al., 2023) with a Nvidia A100. The use of these models allows us to investigate whether our approach holds across different model sizes. In particular, Mistral 7B contains 7 billion parameters, and Mixtral 7x8B 46.7 billion. All of our prompting methods are implemented using the Langchain library.444https://github.com/langchain-ai/langchain We set the decoding temperature to zero and use greedy sampling to make the outputs deterministic. We execute our prompts with in-context learning and provide one demonstration per class. For each experiment, we use a random sample from the training set as demonstrations. The LLM generating the code for code prompts is the same one as the one running the code to generate the final answer. We evaluate each model and prompt in the dev set of each dataset with two random seeds. Since the demonstrations are selected randomly, the seed determines them. The seed that yields the best performance on the dev set is then used for the final evaluation on the test set. We provide more details on the models and the costs in Appendix C and D.
@@ 4.3 Evaluation
| Mistral | Text | 5.05 | |||||
| Code | |||||||
@@ 5 Experiments
We first compare the performance of the two prompting methods — text prompts and code prompts on three LLMs across three datasets (5.1). We then conduct extensive ablation experiments on the dev set of the datasets with GPT 3.5, the best-performing and largest model, to understand the reason behind the performance gain from code prompting. In particular, we study whether code syntax or the implicit text simplification from the code translation is what improves performance (Section 5.2). We also check if the improvement is caused by the models merely being exposed to code within prompts and not necessarily the instances translated to code (Section 5.3). Finally, we show that code prompting is more sample efficient (Section 5.4) when compared to text prompting and that models prompted with code exhibit superior state tracking capabilities (Section 5.5).
@@ 5.1 Code Prompting Improves over Text Prompting
Table 1 shows the model performance on the development and test sets. Code prompts outperform text prompts in the majority of cases on the test set (11 out of 15). This trend holds true across models, with each achieving peak performance through code prompts for most datasets (i.e., GPT-3.5 in 4/5, Mixtral in 3/5, Mistral in 4/5). Notably, code prompts consistently surpass text prompts on BGQA-2 and BGQA-3, the most reasoning-intensive datasets (see Appendix A), for all models. This is particularly evident for GPT-3.5, where gains exceed 18 points. Conversely, the advantage is narrower on CondQA, where the linguistic dimension plays the biggest role (see Appendix A). This suggests that code prompts elicit conditional reasoning abilities and are most suited for reasoning-intensive tasks. Furthermore, in the cases where text prompts are superior, their average gains are only 3.29. In contrast, code prompts lead to significantly larger mean and median gains of 10.48 and 7.75, respectively, in the cases where they are superior. Additionally, an experiment with Phi-2, a small language model, reveals a substantial 15-point performance improvement using code prompts (see Appendix E).
To shed light on the performance gains driven by code prompts, we delve into the confusion matrices (attached in Appendix J) and discover that text prompts in Mistral predict “not enough information” much less than code prompts for BGQA. This is particularly noticeable in BGQA-1, where text prompts do not predict a single “not enough information,” while code prompts do. On the other hand, text prompts in GPT 3.5 and Mixtral overpredict “not enough information” on BGQA and ShARC, leading to a low number of true positives for the conclusive answers. We hypothesize that this model hesitation could stem from the alignment tax (Ouyang et al., 2022) of reinforcement learning from human feedback models. This potential barrier may be alleviated by code prompts because they indicate to the model the variable that answers the question and instruct the model to track the entailment status of variables within the given code.

### v3

@@ Abstract
Reasoning is a fundamental component of language understanding. Recent prompting techniques, such as chain of thought, have consistently improved LLMs’ performance on various reasoning tasks. Nevertheless, there is still little understanding of what triggers reasoning abilities in LLMs in the inference stage. In this paper, we investigate the effect of the input representation on the reasoning abilities of LLMs. We hypothesize that representing natural language tasks as code can enhance specific reasoning abilities such as entity tracking or logical reasoning. To study this, we propose code prompting, a methodology we operationalize as a chain of prompts that transforms a natural language problem into code and directly prompts the LLM using the generated code without resorting to external code execution. We find that code prompting exhibits a high-performance boost for multiple LLMs (up to percentage points on GPT 3.5, on Mixtral, and on Mistral) across multiple conditional reasoning datasets. We then conduct comprehensive experiments to understand how the code representation triggers reasoning abilities and which capabilities are elicited in the underlying models. Our analysis on GPT 3.5 reveals that the code formatting of the input problem is essential for performance improvement. Furthermore, the code representation improves sample efficiency of in-context learning and facilitates state tracking of entities.111Our code and prompts are available at this URL.
@@
Haritz Puerto1, Martin Tutek2††thanks: Work done while author was at TU Darmstadt., Somak Aditya3, Xiaodan Zhu1,4, Iryna Gurevych1 1Ubiquitous Knowledge Processing Lab (UKP Lab), TU Darmstadt and Hessian Center for AI (hessian.AI) 2Technion – IIT, 3IIT Kharagpur, 4Queen’s University https://www.ukp.tu-darmstadt.de
@@ 1 Introduction
Reasoning is a fundamental component of both human and artificial intelligence (AI) and the backbone of many NLP tasks. Recently, intensive studies have been performed on different aspects or types of reasoning such as mathematical reasoning (Patel et al., 2021; Chen et al., 2021b; Cobbe et al., 2021), various kinds of logical reasoning (Liu et al., 2020, 2023a; Sinha et al., 2019), and commonsense-focused reasoning (Madaan et al., 2022; Liu et al., 2022a, b; Wang et al., 2023). Conditional reasoning, a primary yet complex reasoning ability that draws alternative conclusions depending on the fulfillment of certain conditions, remains understudied. These conditions are stated in the text, making the problem self-contained, which allows us to study the semantic inferencing capabilities of the underlying model, i.e., identifying relevant premises and ascertaining the presence of total evidence Nolt et al. (1988); Cabria and Magnini (2014) without the requirement for, and confounding effects of external knowledge. Conditional reasoning is also a fundamental form of reasoning useful in many practical scenarios, such as answering real-world questions about the eligibility for a visa or a loan. Despite the recent introduction of some benchmarks (Saeidi et al., 2018; Sun et al., 2022; Kazemi et al., 2023), conditional reasoning abilities of LLMs remain understudied.
@@
Recently, researchers have analyzed the synergies between LLMs and symbolic interpreters to improve performance on reasoning tasks (Gao et al., 2023; Chen et al., 2023; Lyu et al., 2023). These works transform structured reasoning problems, such as mathematic or symbolic reasoning, into code and run it on an external interpreter. In such a setup, LLMs are mainly focused on natural language representation aspects and planning how to solve the problem, while the actual logical reasoning is offloaded to an external execution module, confounding our understanding of the reasoning In particular, the fundamental questions of what contributes to the reasoning abilities and how reasoning abilities are triggered in LLMs remain open. Nevertheless, pretraining on code is considered an important component that contributes to and explains the improved reasoning ability of LLMs. State-of-the-art LLMs such as GPT 3.5 (Kojima et al., 2022), GPT 4 (OpenAI, 2023), Mixtral (Jiang et al., 2024), and Mistral 7B (Jiang et al., 2023) have been pretrained on both text and code and have demonstrated considerable boosts in many reasoning benchmarks.
In this work, we analyze whether one can elicit improved conditional reasoning abilities in LLMs by merely changing the input format, i.e., from text to code. We constrain our experiments to text+code LLMs to run text and code inputs on the same underlying model. In this way, we can avoid the confounding factor of different pretraining data of specialized text and code LLMs. To understand the benefit of code as an intermediate representation, we devise a chain of prompts, code prompting, that transforms a natural language (NL) task into code and directly prompts the LLM with the generated code. The code contains the logical structure needed to solve the problem, along with the original natural language text as code comments. An illustration is provided in Figure 1. Our contributions are summarized as follows:
We propose a methodology to investigate how the input representation impacts the reasoning abilities of text+code LLMs.
We operationalize such methodology by introducing a chain of prompts that transforms a NL task into code, which is then sent back to the LLM to generate NL answers.
@@
We conduct extensive analysis to understand why code prompts efficiently elicit conditional reasoning abilities, showing that prompting with code yields largely improved variable state tracking.
@@ 2 Background and Related Work
LLM Types.
We categorize LLMs into three types according to their intended use: i) LLMs for natural language text (text-only LLMs), ii) LLMs for coding tasks (code-only LLMs), and iii) LLMs for natural language and coding tasks (text+code LLMs). The intended use of text-only LLMs (Zhang et al., 2022; Touvron et al., 2023a) is to process and generate natural language text such as answers to questions. The intended use of code-only LLMs (Li et al., 2023b; Roziere et al., 2023) is to process and generate code. Lastly, text+code LLMs are equally capable of solving natural language and coding tasks. Examples of this are GPT 3.5 (Kojima et al., 2022), Mixtral (Jiang et al., 2024), and Mistral (Jiang et al., 2023). In this work, we focus on text+code LLMs because of their ability to process two types of input representations interchangeably: natural language text and code. Using such models eliminates the confounding effect of fine-tuning between model variants specialized for only text or code.
Augmenting text with code.
@@
Most works that generate code to solve natural language tasks use an external symbolic interpreter to run the resulting code. Chen et al. (2023) and Gao et al. (2023) showed consistent gains on mathematical problems, symbolic reasoning, and algorithmic problems by using LLMs aided by external code interpreters. Lyu et al. (2023) further report improvements in boolean multi-hop QA, planning, and relational inference. In contrast, Ye et al. (2023) used an external automated theorem prover with declarative code and showed consistent gains w.r.t. imperative code-interpreter-aided LLMs on arithmetic reasoning, logical reasoning, symbolic reasoning, and regex synthesis tasks. Pan et al. (2023) did not use any interpreter and instead created programs composed of multiple subroutines and used smaller specialized models to run them. In this way, they outperform text prompts on text LLMs for fact-checking tasks. Lastly, Li et al. (2023a) runs pieces of code in an LLM to update the program state when the Python interpreter fails due to a code exception and shows performance gains on BIG-Bench Hard (Suzgun et al., 2022). All these works investigate how to best use an external symbolic interpreter to aid an LLM in solving reasoning tasks, i.e., they run code and therefore have a program state with variables and its values. However, we do not employ any external symbolic reasoner, and we do not run code. We investigate the reasoning abilities of LLMs under different input representations (i.e., text and code). Our code prompts are not executed; they are simply read by the LLM and used to generate a natural language answer.
Some works suggest that code LLMs may possess superior reasoning abilities than text LLMs. Madaan et al. (2022) investigate whether code LLMs are superior at structured reasoning than text LLMs. They observe that code LLMs can generate graphs that link commonsense concepts better than text LLMs. Liu et al. (2023b) investigate code prompts in abductive and counterfactual reasoning tasks and report superior results than text prompts on code-davinci (Ouyang et al., 2022), a code LLM. However, code prompts exhibit mixed results on text-davinci-002 (Ouyang et al., 2022), a text LLM. We attribute this to the fact that while this model includes some code in its pretraining corpus, it is not explicitly trained for code generation and, in general, performs poorly on code generation tasks (Chen et al., 2021a). Therefore, the effect of the input representation on the reasoning abilities of text+code LLMs remains unclear. Furthermore, the reasons behind the superior performance of code prompts in code LLMs also remain unclear. In our work, we aim to answer whether code prompts can elicit conditional reasoning abilities in text+code LLMs and the reasons behind this.
@@ 3 Code Prompting
We posit that each LLM encodes a set of capabilities, such as mathematical, logical, or conditional reasoning. However, not all of them are used for every input instance, even if they would be useful. We hypothesize that the input representation plays a pivotal role in eliciting such capabilities. Prior works show that LLMs trained on a combination of text and code exhibit superior reasoning abilities (Kojima et al., 2022; OpenAI, 2023; Jiang et al., 2024, 2023). Therefore, we conjecture that a code representation of a natural language (NL) problem may trigger some of these reasoning abilities encoded in text+code LLMs. More formally, we wonder whether exists some space 222Since the input of LLMs must be strings, must be a set of all possible sentences constructed using some alphabet and grammar. with an associated function that transforms a natural language problem into that space, such that, when prompting an LLM with the representation of in such space yields better results according to some evaluation function .
We fix to the programming language space and define code prompts as prompts that model a natural language problem with code. We also define as a prompt that transforms the NL text into code. code follows the original NL text as much as possible. We use a simple structured code that contains the logical structure needed to solve the problem, along with the original NL text as code comments. In particular, it creates variables for key entities in the question and documents and if blocks for each conditional statement in the documents. Figure 2 exemplifies this transformation and Appendix D provides more details of the code features. Lastly, we define code prompting as , a chain of prompts that i) transform the NL text into code, and ii) use this code to prompt the LLM to generate the answer in natural language. Figure 1 illustrates this pipeline.
It is important to note that the code is not executed per se and therefore, there is no program state. We simply prompt the LLM with the code and ask the LLM to generate a natural language answer based on the content of such code. This setup allows us to investigate the effect of the input representation on text+code LLMs.
@@ 4 Experimental Setup
4.1 Task Setup
We evaluate the conditional reasoning abilities of the LLMs under different prompting methods using the traditional question-answering task. The input is a question and a document, and the model needs to produce the answer, which can be a span of the input document, yes or no. Given that chain-of-thought prompting (CoT, Wei et al. 2022) has been shown to improve the reasoning abilities of LLMs, we instruct the model to generate a CoT before the final answer in all prompting methods. Since we force code prompting to generate a natural language answer, we also force it to generate a natural language CoT.
@@ 4.2 Datasets
We focus on the QA task of CondQA. For BGQA, we focus on the main partition, which includes three subsets BGQA-1, BGQA-2, and BGQA-3, where the number indicates the reasoning hops needed to answer. Lastly, while ShARC encompasses dialogue generation, we aim to evaluate specific capabilities unrelated to conversational flow. Therefore, we isolated the QA pairs from the provided dialogues, resulting in a dataset where the model has to answer yes, no, or not enough information.333In the full task, not enough information would trigger another step in a pipeline to generate a follow-up question. We include more details about the datasets in Appendix B, a formal definition of the prompts in Appendix C, and examples in Appendix Q.
4.3 Models
We perform our study using text+code LLMs because of their ability to process text and code interchangeably. We do not employ code-only LLMs because their intended use does not include solving natural language tasks (Roziere et al., 2023), as required in our case. Similarly, we do not employ text-only LLMs because they cannot generate code. Furthermore, using text+code LLMs also allow us to eliminate the confouding effect of fine-tuning between model variants specialized for only text or code. We corroborate the shortcomings of text-only and code-only LLMs with additional experiments on CodeLLaMA (Roziere et al., 2023) and LLaMA 2 (Touvron et al., 2023b) in Appendix G and H.
We employ OpenAI’s gpt-35-turbo, Mixtral 8x7B (Jiang et al., 2024), and Mistral 7B (Jiang et al., 2023). The use of these models allows us to investigate whether our hypothesis holds across all available sizes of text+code LLMs. We execute our prompts with in-context learning and provide one demonstration per class. More details on the LLM setup are provided in Appendix E.
@@ 4.4 Evaluation
Mistral | Text | 5.10 | |||||
Code |
@@ 5 Experiments
We devise a set of experiments to analyze and quantify whether the code representation of a natural language prompt (i.e., code prompts) elicits conditional reasoning abilities and why. We first compare the performance of the two prompting methods — text prompts and code prompts on three LLMs across three datasets (5.1). We then conduct extensive ablation experiments on the dev set of the datasets with GPT 3.5, the best-performing and largest model, to understand the reason behind the performance gain from code prompting. In particular, we study whether code syntax or the implicit text simplification from the code translation is what improves performance (Section 5.2). We also check if the improvement is caused by the models merely being exposed to code within prompts and not necessarily the instances translated to code (Section 5.3). Furthermore, we show that code prompting is more sample efficient (Section 5.4) when compared to text prompting and that models prompted with code exhibit superior state tracking capabilities (Section 5.5). Lastly, we conduct a human evaluation that confirms the faithfulness of the generated code in Section 6.
@@ 5.1 Code Prompting Improves over Text Prompting
Table 1 shows the model performance on the development and test sets. Code prompts outperform text prompts in the majority of cases on the test set (11 out of 15). This trend holds true across models, with each achieving peak performance through code prompts for most datasets (i.e., GPT-3.5 in 4/5, Mixtral in 3/5, Mistral in 4/5). Notably, code prompts consistently surpass text prompts on BGQA-2 and BGQA-3, the most reasoning-intensive datasets (see Appendix B), for all models. This is particularly evident for GPT-3.5, where gains exceed 18 points. Conversely, the advantage is narrower on CondQA, where the linguistic dimension plays the biggest role (see Appendix B). This suggests that code prompts elicit conditional reasoning abilities and are most suited for reasoning-intensive tasks. Furthermore, in the cases where text prompts are superior, their average gains are only 4.23. In contrast, code prompts lead to significantly larger mean gains of 8.53 in the cases where they are superior. Additionally, an experiment with Phi-2, a small language model, reveals a substantial 15-point performance improvement using code prompts (see Appendix I).
To shed light on the performance gains driven by code prompts, we delve into the confusion matrices (attached in Appendix N) and discover that text prompts in Mistral predict “not enough information” much less than code prompts for BGQA. This is particularly noticeable in BGQA-1, where text prompts do not predict a single “not enough information,” while code prompts do. On the other hand, text prompts in GPT 3.5 and Mixtral overpredict “not enough information” on BGQA and ShARC, leading to a low number of true positives for the conclusive answers. We hypothesize that this model hesitation could stem from the alignment tax (Ouyang et al., 2022) of reinforcement learning from human feedback models. This potential barrier may be alleviated by code prompts because they indicate to the model the variable that answers the question and instruct the model to track the entailment status of variables within the given code.